OpenStudy (anonymous):

evalute the express 6P4

OpenStudy (anonymous):

6 factorial / 4 factorial =6 *5 =30

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

what does P stand for

OpenStudy (anonymous):

6 factorial / 2 factorial =6 *5*4 *3 = 360

OpenStudy (anonymous):

what about for 6P3